OpenStudy (anonymous):

Verify the identity. 4 csc 2(x) = 2 csc^2(x) tan x I work on this for 10 min and couldnt figure out what i did wrong! Help please! :)

OpenStudy (anonymous):

*4csc(2x)* instead of 4csc 2(x)

OpenStudy (anonymous):

\[4\csc2x=\frac{4}{\sin2x}=\frac{4}{2\sin x\cos x}=\frac{2}{\sin x\cos x}\] Try working with that now.

OpenStudy (anonymous):

Thank you! It makes sense now! :D

OpenStudy (anonymous):

\[4 cosec2(x) = 2 cosec^2(x) \tan x \] LHS=4 cosec2(x) \[=\frac{4}{\sin2x}\] \[=\frac{4}{2sinx cosx}=\frac{2}{sinx cosx}\] Multiplying Nr. and Dr. by sinx we find \[\frac{2sinx}{\sin^2x cosx}=2 \frac{1}{\sin^2x }\frac{sinx}{cosx} = 2cosec^2 x \tan x\] =RHS Thus proved